OpenStudy (anonymous):

Explain how to solve the following system of equations. What is the solution to the system? 2x + 2y + z = -5 3x + 4y + 2z = 0 x + 3y + 2z = 1

OpenStudy (mathmale):

What methods of solving systems of equations in 3 variables have you already learned? elimination? matrices? which approach would you prefer to use here?

OpenStudy (anonymous):

I've learned both elimination and matrices. I'd prefer to solve it using elimination though. I'm more comfortable using that to solve problems then matrices.

OpenStudy (mathmale):

Take a look at the 3 equations. Which of the variables, x, y or z, would be easiest to eliminate by addition / subtraction?

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

no wait z!

OpenStudy (mathmale):

Yes, z is a much, much better choice. By what would you have to multiply the first equation so that the coeff. of z is -2?

OpenStudy (anonymous):

-1 right?

OpenStudy (mathmale):

We';ve agreed that we want to eliminate variable z. Look again at the first equation and then decide whether multiplying this equation by -1 or by -2 would be better.

OpenStudy (mathmale):

2x + 2y + z = -5 3x + 4y + 2z = 0 x + 3y + 2z = 1

OpenStudy (anonymous):

oh since the first equation is 1 it would be -2 my b. I was looking at the other two.

OpenStudy (mathmale):

Right. Please multiply the first equation (only) by -2.

OpenStudy (mathmale):

(2x + 2y + z = -5)(-2) 3x + 4y + 2z = 0 x + 3y + 2z = 1

OpenStudy (anonymous):

-4x -4y -2z = 10

OpenStudy (mathmale):

-4x- 4y -2z = 10 Good 3x + 4y + 2z = 0 x + 3y + 2z = 1 Now we're ready to elim. z. First, add the 1st 2 equations together. Cancel the z terms. Next, add the 1st and 3rd eq'ns together. cancel the z terms. You get which 2 equations?

OpenStudy (anonymous):

-1x = 10 -3x -1y = 11 right?

OpenStudy (mathmale):

I haven't actually done the problem. I rather expected you'd get ax + by = c and dx + ey = f But that doesn't mean you're wrong. want to check those calculations once more? combine eq'ns 1 and 2. Next, combine eq'ns 1 and 3 I will do this myself as you do it .

OpenStudy (mathmale):

Mine are the same as yours, except in your very last equation I obtained -3x + y =11. You got -3x - y = 11. Which do you believe is correct?

OpenStudy (anonymous):

I think mine is because you're adding a positive 3y to a negative 4y so there's going to be a -1y left over...right?

OpenStudy (mathmale):

Let's go with yours. -x=10 -3x - y = 11 Are we in agreement?

OpenStudy (anonymous):

Yep

OpenStudy (mathmale):

Pls solve the first equation for x. x=? Next, substitute this numerical value for x in the 2nd equation. What do y ou get?

OpenStudy (mathmale):

Apologies for the long delay. OpenStudy went haywire for 4-5 minutes.

OpenStudy (anonymous):

It's fine. I got x = -10 then I plugged it into the second equation and got y = 19

OpenStudy (mathmale):

Great. Now, how would you find the value of z that satisfies this system of equations?

OpenStudy (anonymous):

plug x and y into the first equation right?

OpenStudy (mathmale):

Any of the original 3 equations would do the trick. You know x and y values, so only z would be unknown. Please do choose the 1st of the original 3 eq'ns. Subst. the known values of x and y and find z.

OpenStudy (anonymous):

ok so x = -10 y = 19 and I got z= -23

OpenStudy (anonymous):

Those are my solutions?

OpenStudy (mathmale):

Substitute your x=-10, y =19 and z = -23 into either the 2nd or the 3rd equation. If the resulting equation is true, your solution set is correct.

OpenStudy (anonymous):

Alright, thank you soo much!

OpenStudy (mathmale):

Thanks for your persistence, Emma. It was a joy to work with you. Sorry for the long delays you experienced.

OpenStudy (anonymous):

Don't worry about it! It was worth it haha